OpenStudy (anonymous):

Simplify completely quantity 2 x squared plus 13 x plus 20 all over x squared minus 5 x minus 36

OpenStudy (anonymous):

the answer that i had got was 2(x+10) ------- x-9 am i right

OpenStudy (anonymous):

@Katt7 @Xmoses1 @varraizr @Nurali @sourwing @Warriorslayer @AngelCriner @LastDayWork

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@bookworm00981 @quixoticideals @Fantalux @Spacelimbus

OpenStudy (fantalux):

Ya i think your right....

OpenStudy (anonymous):

thank you

OpenStudy (fantalux):

No Problem.

OpenStudy (lastdaywork):

@prettigirlrock0804 You made a calculation mistake. Q \[\frac{ 2x^2+13x+20 }{ x^2-5x-36 }\] Your answer \[\frac{ 2(x+10) }{ x-9 }\] Put x = 0 in both the equations and you'll see my point ;)